Find the volume of the solid bounded by

A. x^2 + y^2/ 2 +z = 12 in the region 0 < = x <= 2, 0< =y<=3.

B. f(x,y)=xy in the region x^2 <= y <=x and 0<=x<=1.


1
Expert's answer
2020-01-24T13:29:55-0500

A. V=∫∫(12 -x2-(y2/2))dxdy=∫[12y-x2*y-(y3/6)]30=∫(36-3x2-9/2)dx = (36x-x3-9x/2) |20=72-8-9=55


B. V= ∫∫xydxdy=∫[xy2] xx^2 =0.5*∫(-x5+x3 )dx=((-(x6/12)+(x4/8))|10=1/24=0.416666...